OpenStudy (anonymous):

what is the value of x? 5x+20 x+52

OpenStudy (anonymous):

5x+20 = x+52 5x+20-x=x+52-x 4x+20=52 4x+20-20=52-20 4x=32 x=8

OpenStudy (karatechopper):

x=8

jhonyy9 (jhonyy9):

5x+20 =0 5x= -20 x= -20/5 = -4 x= -4

OpenStudy (karatechopper):

there r 2 equations tht have the same value

jhonyy9 (jhonyy9):

what is x+52 ?

OpenStudy (anonymous):

I think this one needs more information, because 5x + 20 being equal to x + 52 does not make much sense as 5(x+52) is 5x + 260? Don't know been a long time since I did these, but karatechopper's solution is how I would have went about it.